# UniversalCEFR: Enabling Open Multilingual Research on Language Proficiency Assessment

### v1

@@ Abstract
We introduce UniversalCEFR, a large-scale multilingual multidimensional dataset of texts annotated according to the CEFR (Common European Framework of Reference) scale in 13 languages. To enable open research in both automated readability and language proficiency assessment, UniversalCEFR comprises 505,807 CEFR-labeled texts curated from educational and learner-oriented resources, standardized into a unified data format to support consistent processing, analysis, and modeling across tasks and languages. To demonstrate its utility, we conduct benchmark experiments using three modelling paradigms: a) linguistic feature-based classification, b) fine-tuning pre-trained LLMs, and c) descriptor-based prompting of instruction-tuned LLMs. Our results further support using linguistic features and fine-tuning pretrained models in multilingual CEFR level assessment. Overall, UniversalCEFR aims to establish best practices in data distribution in language proficiency research by standardising dataset formats and promoting their accessibility to the global research community.
@@
Table 1: Comparison of existing language learning and language proficiency dataset collections with UniversalCEFR. indicates that only a small subset of the corresponding resource in that repository contains CEFR labels. Among the five repositories, UniversalCEFR is the only non-geo-locked and standardized collection, allowing seamless, unrestricted use for non-commercial research with proper attribution.
@@ 1 Introduction
Recent advances in language proficiency assessment have moved from models relying on hand-crafted linguistic features to large language models (LLMs), which achieve high performance across diverse predictive and generative tasks through post-training techniques such as supervised fine-tuning Devlin et al. (2019); Vaswani et al. (2017) or instruction tuning Wei et al. (2022). This form of task generalization enables complex linguistic pattern (e.g., features that make a text complex) modelling within unified frameworks for assessing language proficiency on standardized scales like CEFR. Moreover, they can also be extended to low-resourced languages, potentially improving automatic assessment through techniques such as cross-lingual transfer He and Li (2024); Imperial and Kochmar (2023a, b); Vajjala and Rama (2018).
@@ 2 Background
Automatic CEFR Assessment.
@@
In parallel with the rise of benchmarking studies for LLMs, similar efforts are growing in the CEFR-based language proficiency community. Two works in this direction include Naous et al. (2024), which introduced ReadMe++, a multilingual, multidomain dataset for sentence-level readability assessment on a CEFR scale covering five languages, while the iRead4Skills Project by Pintard et al. (2024) released a collection of written texts in French, Portuguese, and Spanish across multiple genres and levels patterned to CEFR.
@@ 3 The UniversalCEFR Dataset
To support multilingual language proficiency research, we introduce UniversalCEFR, a large-scale initiative that curates and standardizes open human-annotated CEFR-labeled corpora. In this section, we outline the design principles behind dataset’s development, the data collection and standardization pipeline, key dataset statistics, and a linguistic feature analysis that supports downstream modeling.
@@ 3.1 Design Principles
From its conceptualization and planning, the UniversalCEFR initiative involved close collaboration among 20 researchers in language proficiency assessment, NLP, and education from 13 institutions across nine countries (UK, Canada, USA, Germany, Sweden, UAE, Spain, Belgium, and Portugal). [^7] As CEFR is a European framework, most active researchers in the field are based in Europe. They all played a key role in defining the standardization protocol, designing evaluation experiments, and discussing future research directions. These collaborative decisions are discussed in more detail in the following sections.
@@ 3.2 Data Collection
To ensure interoperability, transformation, and machine readability, we standardized the collected datasets by preprocessing their varied source formats into a unified structure. We adopted JSON as the per-instance format and defined eight metadata fields considered essential for each CEFR-labeled text. These fields include the source dataset, language, granularity (document, paragraph, sentence, discourse), production category (learner or reference), and license. Full descriptions and predetermined values used for each field are provided in Table 15. The final standardized dataset was uploaded to a HuggingFace Dataset repository, [^8] https://github.com/huggingface/datasets and will be made publicly available after the review period. A key challenge was the lack of a unified format across the language proficiency community. Source corpora came in various formats, including plain text (e.g., csv, tsv, txt), spreadsheets (e.g., XLSX, XLS), markup (e.g., XML), and PDFs requiring manual extraction. This challenge further motivates the need for unified data aggregation initiatives that UniversalCEFR aims to help establish.
@@ 3.3 Dataset Statistics
The final UniversalCEFR collection comprises 505,807 CEFR-labeled texts across 13 languages and 4 scripts (Latin, Arabic, Devanagari, and Cyrillic). Table 3.2 shows the overall dataset size and its splits.We identified 11,316 instances with invalid or out-of-scope labels (e.g., NA, A+, B) outside the six recognized CEFR labels (A1-C2) and duplicates, which were removed before splitting UniversalCEFR into train, dev, and test. For the test, we set a cap of 200 instances per language and per granularity level. Additional dataset statistics can be found in Appendix A.
3.4 Linguistic Feature Analysis
To explore the relationship between CEFR levels and linguistic variation in UniversalCEFR, we conducted a Spearman correlation analysis. Length-based measures, such as characters and syllables per sentence, showed the strongest correlation with CEFR levels across all languages. However, this varied significantly across languages: Czech, Estonian, and Italian showed a high number of correlated features, whereas Welsh and German showed very few or none, indicating a limited alignment between the current feature set and CEFR levels in those languages.
In an additional point-biserial correlation analysis by CEFR level (i.e., one level versus all others), most features showed only weak correlations, suggesting limited discriminative power when isolating individual CEFR bands. However, we found an interesting dynamic in some features, e.g., the word length in characters was negatively correlated with A1, neutral with A2, and positive with B1 and higher levels. Hence, the directionality of several features suggests dynamic usage patterns across CEFR bands, even if the correlation strengths remain modest. Further results and detailed analyses are provided in Appendix E.3.
Model & Setup EN ES DE NL CS IT FR ET PT AR HI RU CY Avg
Baseline
Gemma1-7B (English)
Gemma3-12B (Multi)
EuroLLM-9B (Multi)
fine-tuned Models
Feature-Based Models
4 CEFR Level Classification
4.1 Feature-Based Models
We evaluated two widely-used classification models from Scikit-Learn Pedregosa et al. (2011): Random Forest (RandForest) and Logistic Regression (LogRegr). Both models were trained on the linguistic features described in Section 3.4, using Scikit-Learn’s default hyperparameter settings. We experimented with two feature configurations: one using all 100 features (AllFeats) and another using an automatically selected subset of top-performing features across all languages (TopFeats). Appendix E.1 and E.2 detail the linguistic feature information for both setups.
4.2 Fine-tuned Models
4.3 Descriptor-Based Prompting
  • Base. Generic prompting with no CEFR level descriptors as context.
All CEFR descriptors were retrieved from the official CEFR website. Prompt templates and hyperparameter values for each setup are detailed in the Table 18 and Appendix I.
4.4 Evaluation Metrics
We use weighted F1 score as primary evaluation metric across all experiments. This accounts for the class imbalance in CEFR level distribution and granularity across language subsets in UniversalCEFR-test. Using accuracy in the experiments would produce misleading performance in favor of any majority class.
5 Results
5.1 Model-Based Performance Comparison
Table 3 shows that, in terms of overall average performance across languages, the fine-tuned setup with ModernBERT, EuroBERT, and XLM-R achieved the highest weighted F1 score range ($\approx$60%-62.8%) outperforming feature-based models ($\approx$47%-58%) and prompting ($\approx$23%-43%). Among the LLM-based approaches—prompting and fine-tuning—models trained on broader multilingual corpora generally performed better. For instance, XLM-R, which supports 100 languages, was the top performer, followed by EuroBERT (15 languages) and ModernBERT (English-only). A similar trend was observed in prompting: Gemma 3, trained on 140+ languages, outperformed EuroLLM (15 languages) and the English-centric Gemma 1, achieving the best prompting score of 43.2. These findings are consistent with previous work (Naous et al., 2024; Shardlow et al., 2024; Colla et al., 2023; Yuan and Strohmaier, 2021), reinforcing the usefulness of multilingual models for language proficiency assessment tasks. One limitation of our experimental setup, however, is that we did not include language-specific pre-trained models for languages other than English, which may have further improved performance for low- and mid-resource languages.
Model Sent Para Doc All
Gemma1 19.41 42.74 30.81 33.63
Gemma3 38.71 43.12 39.62 42.33
XLM-R 62.67 66.38 71.12 65.92
RandForest-All 56.88 62.77 64.58 61.38
RandForest-Top 53.89 62.98 64.94 60.50
5.2 Granularity-Level Comparison
Table 4 highlights clear performance differences across text granularities (sentence, paragraph, and document) for all models, but more prominently for the Gemma models under prompting. Gemma 1, in particular, tends to over-predict lower CEFR levels (A1-B1) on sentence-level data, whereas its predictions on document-level subsets are more evenly distributed and better aligned with ground truth distributions. This suggests that prompt-based methods may require longer texts to make more accurate predictions, unlike models trained or fine-tuned on the respective datasets. For other models, such as XLM-R, Random Forest, performance show better results with document ($\approx$64%-71%) and paragraph-level data ($\approx$62%-66%) than sentence-level data ($\approx$53%-62%), which was shown as a more difficult task in previous work on readability Dell’Orletta et al. (2011); Vajjala and Meurers (2014). Regarding language-specific differences, among English, German, and Welsh, the best performance is seen with the paragraph level dataset for English, the document level dataset for German, and the sentence level dataset for Welsh and French with the fine-tuned XLM-R model. Similar variations can be observed for other languages with more than one level of granularity (see Table 19). No single granularity or model gives a consistently better performance across all tested languages. These results are likely due to the distribution of excerpts across granularity levels in each language (see Table 7 in Appendix A).
5.3 Learner-Reference Comparison
Four languages in UniversalCEFR contain both learner and reference texts: Arabic, German, English, and Spanish. Table 5 reports the average weighted F1 performance difference between the two categories across the four languages. For German, performance is comparable between learner and reference texts ($\approx$71–74%). In contrast, English and Spanish show higher performance on learner texts (83% and 98%) than on reference texts (58% and 42%, respectively). Arabic displays the opposite trend: results on reference texts (54%) are much higher than those of learner texts, where the best results was obtained by Gemma 3 (41%). One possible explanation is that Gemma 3 may have been exposed to more Arabic content in its pre- and post-training phases.
Language Learner Reference
AR 41.92 54.69
DE 71.14 74.39
EN 83.41 58.24
ES 97.99 42.72
Table 5: Average performances of the best models on learner text versus reference text across languages. indicates performance with Gemma3, and the rest are the performance of the XLM-R model. Only these four languages have both learner and reference texts.
6 Discussion
Critical Reflections of Current Practices.
The multiregional and multidisciplinary effort to curate UniversalCEFR provided us with a comprehensive view of current inconsistencies and critical gaps in building CEFR-labeled language proficiency assessment corpora. On examining the annotation practices, there appears to be no standard method for conducting expert annotations, including inconsistent use of inter-annotator agreement metrics and unclear guidelines on the number of annotators required to achieve reliable agreement. This is reflected in the UniversalCEFR dataset, where nearly half of the corpora lack information on the annotators involved and their agreement scores. We posit that this may be due to diverse judgment of what constitutes high-quality data that does not require further human annotations.
In terms of language coverage, UniversalCEFR includes nine (EN, ES, DE, NL, CS, IT, FR, ET, PT) of the 24 recognized European languages. [^9] https://european-union.europa.eu/principles-countries-history/languages_en As a result, researchers working on these nine languages now have access to open, standardized data for CEFR-based language proficiency assessment. The remaining 15 languages represent valuable opportunities for future expansion through collaborative efforts. While our open data and standardization initiative is a step towards addressing current challenges in interoperability and accessibility of resources, similar parallel efforts are needed in areas such as annotation and evaluation practices to ensure sustained progress in the language proficiency assessment community.
Need for Pro-Research Data Sharing Policies.
Linguistic Features and Fine-tuning Still Matter.
While recent advances in LLMs keep transforming NLP research, our multidimensional experiments in Section 5 reaffirm the continued value of linguistic features for traditional ML classifiers and fine-tuning pretrained models in language proficiency assessment. We observe common patterns where higher distribution and instance count lead to better results using these two setups (see performances on Spanish, English, and German subsets in Table 3) over prompting with CEFR descriptors. Moreover, using linguistic features in language proficiency assessment allows deeper analysis of language interactions with variables such as complexity. For a more detailed discussion, Appendix C provides language-specific analyses across model predictions and possible factors influencing performance, including potential effects of scale, and language exposure from multilingual models.
7 Conclusion and Future Directions
In this work, we introduced UniversalCEFR, a large-scale, open, multilingual, multidimensional dataset comprising 505,807 CEFR-annotated texts across 13 languages developed through global collaboration. Our findings through diverse model experiments with CEFR level prediction strengthen support for the utility of linguistic features and fine-tuning multilingual models in language proficiency assessment. Similarly, our critical analysis of the current data and resource-building practices emphasized the need for similar initiatives from the community, and pro-research data sharing policies in the advent of generative AI to remove barriers to accessibility without compromising data privacy and intellectual property.
Beyond its data and technical contributions, UniversalCEFR also carries broader sociolinguistic significance. UniversalCEFR addresses the growing linguistic inequality in modern AI development through focusing on under-represented languages alongside English. We hope this initiative can lead to more responsible AI development that actively resists the growing linguistic centralization around English in global AI research—a modern Matthew effect Merton (1988)—where well-resourced languages receive disproportionate technological attention while smaller languages (like Czech or Welsh) are left behind Masciolini et al. (2025). The UniversalCEFR is a strong step towards mitigating the Matthew effect.
Limitations
We discuss several limitations of our work for UniversalCEFR and how researchers can consider these directions to develop the resource further.
Language Availability and Dependency.
Due to the nature of UniversalCEFR being a collection of open-sourced, publicly accessible CEFR data, its growth depends heavily on how the community will move forward and continuously release artifacts, including CEFR-annotated corpora for reproducibility and wider access for research purposes. We also give credit to the efforts of researchers who work on multi-framework adoption, where CEFR descriptors and bands are overlapped with languages not within Europe (such as Hindi Naous et al. (2024) and Arabic Habash and Palfreyman (2022)), and continue to open-source the annotated data.
Natural Data Disparity Across Languages.
From the statistics presented in Tables 6 and 7 for UniversalCEFR, it is expected that not all languages have the exact same distribution of data across dimensions. Research efforts on CEFR with the English language remain the portion with the largest open-sourced dataset, being an international language. Moreover, languages such as Welsh do not have any openly accessible CEFR-labeled data at all, to which UniversalCEFR is set to be the first contributor for A1 and A2, covering sentence-, paragraph-, document-, and dialogue-level formats.
Modalities Beyond Texts.
Ethics Statement
In the context of the EU AI Act, the use of AI systems for educational purposes is classified under high risk, especially those that are intended to "to evaluate learning outcomes, including when those outcomes are used to steer the learning process of natural persons in educational and vocational training institutions at all levels." European Parliament and Council (2024). Thus, AI systems that will be released in the market with these goals are required to comply with obligations for high-risk systems, including data governance with high-quality, representative datasets. As a form of contribution towards meeting these requirements, the UniversalCEFR is an initiative that will allow researchers and developers access to diverse, multilingual, multidimensional CEFR-labeled texts which can be used for designing systems that are representative, explainable, and fair.
Acknowledgments
Appendix A Full Data Statistics
Tables 7, 9, 11and 13 report the quantity of CEFR-labeled texts across granularity levels per language, and Tables 6, 8, 10 and 12 reflect their counterpart in terms of CEFR level coverage. In forming the test split, we randomly sampled CEFR-labeled text instances per language per granularity level, while setting a cap of 200. This allows us to have a sizeable representation of UniversalCEFR while maintaining efficiency for inference with LLMs. In total, we have 4,465 CEFR-labeled instances for UniversalCEFR-test, which is comparable to the general sizes of benchmark test sets from previous works related to language proficiency Naous et al. (2024); Zhang et al. (2024); Imperial and Tayyar Madabushi (2024). For the train and dev sets for fine-tuning and feature-based classification, we split the full subset (minus the test set) into a 90%-10% partition, respectively.

### v2

@@ Abstract
We introduce UniversalCEFR, a large-scale multilingual and multidimensional dataset of texts annotated with CEFR (Common European Framework of Reference) levels in 13 languages. To enable open research in automated readability and language proficiency assessment, UniversalCEFR comprises 505,807 CEFR-labeled texts curated from educational and learner-oriented resources, standardized into a unified data format to support consistent processing, analysis, and modelling across tasks and languages. To demonstrate its utility, we conduct benchmarking experiments using three modelling paradigms: a) linguistic feature-based classification, b) fine-tuning pre-trained LLMs, and c) descriptor-based prompting of instruction-tuned LLMs. Our results support using linguistic features and fine-tuning pretrained models in multilingual CEFR level assessment. Overall, UniversalCEFR aims to establish best practices in data distribution for language proficiency research by standardising dataset formats, and promoting their accessibility to the global research community.
@@
Table 1: Comparison of existing language learning and language proficiency dataset collections with UniversalCEFR. indicates that only a subset of the corresponding resource in that repository contains CEFR labels. Among the five repositories, UniversalCEFR is the only non-geo-locked and standardized collection, allowing seamless, unrestricted use for non-commercial research with proper attribution.
@@ 1 Introduction
Recent advances in language proficiency assessment have moved from models relying on hand-crafted linguistic features to large language models (LLMs), which achieve high performance across diverse predictive and generative tasks through post-training techniques such as supervised fine-tuning Devlin et al. (2019); Vaswani et al. (2017) or instruction tuning Wei et al. (2022). This form of task generalization enables complex linguistic pattern (e.g., features that make a text complex) modelling within unified frameworks for assessing language proficiency on standardized scales like CEFR. Moreover, they can also be extended to low-resource languages, potentially improving automatic assessment through techniques such as cross-lingual transfer He and Li (2024); Imperial and Kochmar (2023a, b); Vajjala and Rama (2018).
@@ 2 Background
CEFR Assessment and Standardization.
@@
In parallel with the rise of benchmarking studies for LLMs, similar efforts are growing in the CEFR-based language proficiency community. Two works in this direction include Naous et al. (2024), which introduced ReadMe++, a multilingual, multidomain dataset for sentence-level readability assessment on a CEFR scale covering five languages, while the iRead4Skills Project by Pintard et al. (2024) released a collection of written texts in French, Portuguese, and Spanish across multiple genres and levels patterned to CEFR. Likewise, in data collection standardization, CLARIN released the Core Metadata Schema for Learner Corpora (LC-meta), which aims to provide a structured method with a specific emphasis on capturing metadata of collected learner texts, focusing on learner background, context, and individual differences Paquot et al. (2024).
@@ 3 The UniversalCEFR Dataset
To support multilingual language proficiency research, we introduce UniversalCEFR, a large-scale initiative that curates and standardizes open human-annotated CEFR-labeled corpora. Unifying diverse resources under a consistent format enables reproducible and scalable research across linguistics, NLP, and education. In this section, we outline the dataset’s design principles, detail the data collection and standardization pipeline, provide key statistics, and present a linguistic feature analysis that supports downstream modelling.
@@ 3.1 Design Principles
From its conceptualization and planning, the UniversalCEFR initiative involved close collaboration among 20 researchers in language proficiency assessment, NLP, and education from 13 institutions across nine countries (UK, Canada, USA, Germany, Sweden, UAE, Spain, Belgium, and Portugal). [^7] As CEFR is a European framework, most active researchers in the field are based in Europe. They all played a key role in defining the standardization protocol, designing evaluation experiments, and discussing future research directions. These collaborative decisions are detailed in the following sections.
@@ 3.2 Data Collection
To ensure interoperability, transformation, and machine readability, we standardized the collected datasets by preprocessing their varied source formats into a unified structure. We adopted JSON as the per-instance format and defined eight metadata fields considered essential for each CEFR-labeled text. These fields include the source dataset, language, granularity (document, paragraph, sentence, discourse), production category (learner or reference), and license. Full descriptions and predetermined values used for each field are provided in Table 15. The final standardized dataset is available from HuggingFace Dataset repository. [^8] https://huggingface.co/UniversalCEFR/datasets A key challenge was the lack of a unified format across the language proficiency community. Source corpora came in various formats, including plain text (e.g., csv, tsv, txt), spreadsheets (e.g., XLSX, XLS), markup (e.g., XML), and PDFs requiring manual extraction. This challenge further motivates the need for unified data aggregation initiatives that UniversalCEFR aims to help establish.
@@ 3.3 Dataset Statistics
The final UniversalCEFR collection comprises 505,807 CEFR-labeled texts across 13 languages and 4 scripts (Latin, Arabic, Devanagari, and Cyrillic). Tables 2 and  3 show the overall dataset size, its splits and breakdown per CEFR level per language. We identified 11,316 instances with invalid or out-of-scope labels (e.g., NA, A+, B) outside the six recognized CEFR labels (A1–C2) and duplicates, which were removed before splitting UniversalCEFR into train, dev, and test. For the test, we set a cap of 200 instances per language and per granularity level. Additional dataset statistics can be found in Appendix A.
@@
4 Linguistic Feature Analysis
4.1 Correlation Across All Languages
Figure 2: Highly correlated linguistic features occurring in at least three languages. Blue features lean towards positive correlation, while red features denote negative correlation. For brevity, these top features are those lying at the extreme ends of the correlation spectrum.
4.2 Correlation By CEFR Level
To assess the consistency of feature relevance across languages, we examined the number of features with significant correlations ($p<0.05$) with CEFR levels per language as visualized in Figure 2. The results revealed notable variations. Languages such as Czech (cs), Estonian (et), and Italian (it) showed a high number of relevant features, suggesting strong alignment between the selected linguistic features and CEFR progression in these languages. English (en), Spanish (es), French (fr), Hindi (hi), and Russian (ru) showed moderate coverage, with a reasonable number of features exceeding the 0.3 correlation threshold. In contrast, Arabic (ar), Dutch (nl), and Portuguese (pt) exhibited weak coverage, while Welsh (cy) and German (de) had very few or no features with relevant correlations, indicating a limited match between the current feature set and CEFR levels for those languages. Furthermore, a few features are only relevant for a few languages, e.g., the translative case for only Estonian, negative verb polarity for only Czech, or genitive case for only Czech, Estonian, and Russian. This variability highlights the influence of language-specific properties on the effectiveness of general feature-based models for proficiency prediction.
Model & Setup EN ES DE NL CS IT FR ET PT AR HI RU CY Avg
[…]
5 CEFR Level Classification
5.1 Feature-Based Models
We evaluated two widely-used classification models from Scikit-Learn Pedregosa et al. (2011): Random Forest (RandForest) and Logistic Regression (LogRegr). Both models were trained on the linguistic features described in Section 4, using Scikit-Learn’s default hyperparameter settings. We experimented with two feature configurations: one using all 100 features (AllFeats) and another using an automatically selected subset of top-performing features across all languages (TopFeats). Appendices E.1 and E.2 detail the linguistic feature information for both setups.
5.2 Fine-tuned Models
5.3 Descriptor-Based Prompting
  • Base. Generic prompting with no CEFR level descriptors as context.
All CEFR descriptors were retrieved from the official CEFR website. Prompt templates and hyperparameter values for each setup are detailed in Table 18 and Appendix I.
5.4 Evaluation Metrics
We use weighted F1 as the primary evaluation metric across all experiments. This accounts for the class imbalance in CEFR level distribution and granularity across language subsets in UniversalCEFR-test. Using accuracy in the experiments would produce misleading performance in favor of any majority class.
6 Results
6.1 Model-Based Performance Comparison
Table 4.2 shows that, in terms of overall average performance across languages, the fine-tuned setup with ModernBERT, EuroBERT, and XLM-R achieved the highest weighted F1 score range ($\approx$60%-62.8%) outperforming feature-based models ($\approx$47%-58%) and prompting ($\approx$23%-43%). Among the LLM-based approaches—prompting and fine-tuning—models trained on broader multilingual corpora generally performed better. For instance, XLM-R, which supports 100 languages, was the top performer, followed by EuroBERT (15 languages) and ModernBERT (English-only). A similar trend was observed in prompting: Gemma 3, trained on 140+ languages, outperformed EuroLLM (15 languages) and the English-centric Gemma 1, achieving the best prompting score of 43.2. These findings are consistent with previous work (Naous et al., 2024; Shardlow et al., 2024; Colla et al., 2023; Yuan and Strohmaier, 2021), reinforcing the usefulness of multilingual models for language proficiency assessment tasks. One limitation of our experimental setup, however, is that we did not include language-specific pre-trained models for languages other than English, which may have further improved performance for low- and mid-resource languages.
Model Sent Para Doc All
Gemma1 19.41 42.74 30.81 33.63
Gemma3 38.71 43.12 39.62 42.33
XLM-R 62.67 66.38 71.12 65.92
RandForest-All 56.88 62.77 64.58 61.38
RandForest-Top 53.89 62.98 64.94 60.50
6.2 Granularity-Level Comparison
Table 5 highlights clear performance differences across text granularities (sentence, paragraph, and document) for all models, but more prominently for the Gemma models under prompting. Gemma 1, in particular, tends to over-predict lower CEFR levels (A1–B1) on sentence-level data, whereas its predictions on document-level subsets are more evenly distributed and better aligned with ground truth distributions. This suggests that prompt-based methods may require longer texts to make more accurate predictions, unlike models trained or fine-tuned on the respective datasets. Other models, such as XLM-R and Random Forest, show better results on document ($\approx$64%-71%) and paragraph-level data ($\approx$62%-66%) than sentence-level data ($\approx$53%-62%), which was shown to be a more difficult task in previous work on readability Dell’Orletta et al. (2011); Vajjala and Meurers (2014). Regarding language-specific differences, among English, German, and Welsh, the best performance is seen with the paragraph-level dataset for English, the document-level dataset for German, and the sentence-level dataset for Welsh and French with the fine-tuned XLM-R model. Similar variations can be observed for other languages with more than one level of granularity (see Table 19). No single granularity or model shows consistently better performance across all tested languages. These results are likely due to the distribution of excerpts across granularity levels in each language (see Table 7 in Appendix A).
6.3 Learner-Reference Comparison
Four languages in UniversalCEFR contain both learner and reference texts: Arabic, German, English, and Spanish. Table 6 reports the average weighted F1 performance difference between the two categories across the four languages. For German, performance is comparable between learner and reference texts ($\approx$71–74%). In contrast, English and Spanish show higher performance on learner texts (83% and 98%) than on reference texts (58% and 42%, respectively). Arabic displays the opposite trend: results on reference texts (54%) are much higher than those of learner texts, where the best results were obtained by Gemma 3 (41%). One possible explanation is that Gemma 3 may have been exposed to more Arabic content in its pre- and post-training phases.
Language Learner Reference
AR 41.92 54.69
DE 71.14 74.39
EN 83.41 58.24
ES 97.99 42.72
Table 6: Average performances of the best models on learner text versus reference text across languages. indicates performance with Gemma 3, and the rest refer to performance of the XLM-R model. Only these four languages have both learner and reference texts.
7 Discussion
Critical Reflections of Current Practices.
The multiregional and multidisciplinary effort behind UniversalCEFR exposed significant inconsistencies and critical gaps in building CEFR-labeled language proficiency assessment corpora. Upon examination of annotation practices, there appears to be no standard method for conducting expert annotations, including inconsistent use of inter-annotator agreement metrics and unclear guidelines on the number of annotators required to achieve reliable agreement. This is reflected in the UniversalCEFR dataset itself, where nearly half of the corpora lack information on the annotators involved and their agreement scores. We posit that this may be due to diverse judgments of what constitutes high-quality data that does not require further human annotations.
In terms of language coverage, UniversalCEFR includes nine (EN, ES, DE, NL, CS, IT, FR, ET, PT) of the 24 recognized European languages. As a result, researchers working on these nine languages now have access to open, standardized data for CEFR-based language proficiency assessment. The remaining 15 languages represent valuable opportunities for future expansion through collaborative efforts. While our open data and standardization initiative is a step towards addressing current challenges in interoperability and accessibility of resources, similar parallel efforts are needed in areas such as annotation and evaluation practices to ensure sustained progress in the language proficiency assessment community.
Need for Pro-Research Data Sharing Policies.
Linguistic Features and Fine-tuning Still Matter.
While recent advances in LLMs keep transforming NLP research, our multilingual and multidimensional experiments in Section 6 reaffirm the continued value of linguistic features for traditional ML classifiers and fine-tuning pre-trained models in language proficiency assessment. We observe common patterns where higher distribution and instance count lead to better results using these two setups (see performances on Spanish, English, and German subsets in Table 4.2) over prompting with CEFR descriptors. Moreover, using linguistic features in language proficiency assessment allows deeper analysis of language interactions with variables such as complexity, as seen in Appendix C. Given these insights, we encourage further efforts in the expansion of existing but low-resource language datasets with CEFR labels, as well as the exploration of features to better model morphologically-rich languages (e.g., Estonian and Portuguese). Together, these recommendations bridge current observed model failures to practical approaches in improving multilingual CEFR proficiency assessment.
8 Conclusion and Future Directions
In this work, we introduced UniversalCEFR, a large-scale, open, multilingual, multidimensional dataset comprising 505,807 CEFR-annotated texts across 13 languages developed through global collaboration. Our findings from diverse model experiments with CEFR level prediction provide strong support for the utility of linguistic features and fine-tuning multilingual models in language proficiency assessment. Similarly, our critical analysis of the current data and resource-building practices emphasized the need for similar initiatives from the community, and pro-research data sharing policies in the advent of generative AI to remove barriers to accessibility without compromising data privacy and intellectual property.
Beyond its data and technical contributions, UniversalCEFR also carries broader sociolinguistic significance. UniversalCEFR addresses the growing linguistic inequality in modern AI development by focusing on underrepresented languages alongside English. We hope this initiative can lead to more responsible AI development that actively resists the growing linguistic centralization around English in global AI research—a modern Matthew effect Merton (1988)—where well-resourced languages receive disproportionate technological attention while smaller languages (like Czech or Welsh) are left behind Masciolini et al. (2025). The UniversalCEFR is a strong step towards mitigating the Matthew effect in language proficiency assessment research.
Limitations
We discuss several limitations of our work on UniversalCEFR and how researchers can consider these directions to develop the resource further.
Natural Data Disparity in Experiments.
From the statistics presented in Tables 3 and 7 for UniversalCEFR, it is expected that not all languages have the exact same distribution of data across dimensions, including formats (sentence-, paragraph-, document-, and dialogue-level) and category (reference and learner texts). Hence, our main experiments in Table 4.2 combined these variables to provide a unified performance comparison. We note that while this offers a broad overview of the three evaluation paradigms (prompting, fine-tuning, and linguistic features) across languages, future work should include dedicated modeling and evaluation by text category, which may warrant a more focused, in-depth study we assign for future work.
Language Availability and Dependency.
Due to the nature of UniversalCEFR being a standardized collection of open-sourced, publicly accessible CEFR data, its growth depends heavily on how the community will move forward and continuously release artifacts, including CEFR-annotated corpora for reproducibility and wider access for research purposes. We also acknowledge the efforts of researchers who work on multi-framework adoption, where CEFR descriptors and bands are overlapped with languages not within Europe (such as Hindi Naous et al. (2024) and Arabic Habash and Palfreyman (2022)), and continue to open-source the annotated data.
Modalities Beyond Texts.
Beyond Typical Benchmarking
The rigor of analysis in this paper is not meant to be treated as a typical benchmark study, similar to recent trends in NLP papers, where the goal is to evaluate as many LLMs as possible. In this paper, we provide deeper insights into language complexities and intricacies that affect model performance in CEFR level classification across various dimensions of language, granularity, and format. Thus, within our compute budget, we carefully handpicked state-of-the-art LLMs that are worth exploring based on their properties (e.g., English-centric against massively multilingual, or linguistic features against fine-tuning and prompting). We leave the evaluation on larger, more advanced LLMs, as well as explorations in other directions to improve CEFR level classification, such as the use of high-quality synthetic datasets, for future work.
Ethics Statement
In the context of the EU AI Act, the use of AI systems for educational purposes, especially those that are intended to "to evaluate learning outcomes, including when those outcomes are used to steer the learning process of natural persons in educational and vocational training institutions at all levels" European Parliament and Council (2024), is classified under high risk. Thus, AI systems that will be released in the market with these goals are required to comply with obligations for high-risk systems, including data governance with high-quality, representative datasets. As a form of contribution towards meeting these requirements, the UniversalCEFR is an initiative that will allow researchers and developers access to diverse, multilingual, multidimensional CEFR-labeled texts which can be used for designing systems that are representative, explainable, and fair.
Acknowledgments
Appendix A Full Data Statistics
Tables 7, 9, 11and 13 report the quantity of CEFR-labeled texts across granularity levels per language, and Tables 3, 8, 10 and 12 reflect their counterparts in terms of CEFR level coverage. In forming the test split, we randomly sampled CEFR-labeled text instances per language per granularity level, while setting a cap of 200. This allows us to have a sizeable representation of UniversalCEFR while maintaining efficiency for inference with LLMs. In total, we have 4,465 CEFR-labeled instances for UniversalCEFR-test, which is comparable to the general sizes of benchmark test sets from previous works related to language proficiency Naous et al. (2024); Zhang et al. (2024); Imperial and Tayyar Madabushi (2024). For the train and dev sets for fine-tuning and feature-based classification, we split the full subset (minus the test set) into a 90%-10% partition, respectively.